# Harnessing PDF Data for Improving Japanese Large Multimodal Models

### v1

@@ Abstract
Large Multimodal Models (LMMs) have demonstrated strong performance in English, but their effectiveness in Japanese remains limited due to the lack of high-quality training data. Current Japanese LMMs often rely on translated English datasets, restricting their ability to capture Japan-specific cultural knowledge. To address this, we explore the potential of Japanese PDF data as a training resource, an area that remains largely underutilized. We introduce a fully automated pipeline that leverages pretrained models to extract image-text pairs from PDFs through layout analysis, OCR, and vision-language pairing, removing the need for manual annotation. Additionally, we construct instruction data from extracted image-text pairs to enrich the training data. To evaluate the effectiveness of PDF-derived data, we train Japanese LMMs and assess their performance on the Japanese LMM Benchmark. Our results demonstrate substantial improvements, with performance gains ranging from 3.9% to 13.8% on Heron-Bench. Further analysis highlights the impact of PDF-derived data on various factors, such as model size and language models, reinforcing its value as a multimodal resource for Japanese LMMs. We plan to make the source code and data publicly available upon acceptance.
@@
Jeonghun Baek Akiko Aizawa Kiyoharu Aizawa The University of Tokyo  National Institute of Informatics baek@hal.t.u-tokyo.ac.jp
@@ 1 Introduction
We train Japanese LMMs with the PDF-derived data and evaluate them on the Japanese LMM Benchmark. Our results show that PDF-derived data significantly improves performance, with gains of 3.9% to 13.8% on Heron-Bench Inoue et al. (2024). Furthermore, we conduct additional experiments to provide various insights into PDF-derived data’s effectiveness. Our main contributions are as follows:
@@
  • We demonstrate that PDF-derived data can significantly improve Japanese LMM performance, achieving 3.9% to 13.8% gains on Heron-Bench.
@@ 2.2 Japanese LMM
Some models, such as Qwen-VL Bai et al. (2023), achieve strong performance in Japanese without using a Japanese LLM, instead leveraging a multilingual LLM. VILA-jp Sasagawa et al. (2024) has further improved performance by utilizing interleaved data. However, no existing work has leveraged PDF data to enhance Japanese LMMs. To achieve higher performance, we utilize PDF data in our approach.
@@ 4.1 Training Procedure
Training Data.  Table 1 presents the details of the training data. In Stage 1 (Pretraining) and Stage 2 (Instruction Tuning), we use the Japanese-translated version of the original LLaVA training data Inoue et al. (2024), translated using DeepL DeepL . Specifically, Stage 1 uses 558K samples from LLaVA-Pretrain-JA Motors (2024a), Stage 2 uses 620K samples from LLaVA-v1.5-Instruct-620K-JA Motors (2024b), and Stage 3 uses 362K PDF-derived samples created in §3.
@@ 4.2 LLM Selection
Japanese LLM.  We train LLaVA1.5 using three well-known Llama3-8B-based Japanese LLMs: Suzume (suzume-llama-3-8B-japanese Devine (2024)), ELYZA (Llama-3-ELYZA-JP-8B Hirakawa et al. (2024)), and Swallow (Llama-3-Swallow-8B-Instruct-v0.1 Okazaki et al. (2024))
@@
Dataset #Questions #Images JA-LLaVA-Bench (COCO) 90 30 JA-LLaVA-Bench (Wild) 60 24 Heron-Bench 102 21
Table 2: Details of Japanese LMM evaluation sets.
JA-LLaVA-Bench (COCO) JA-LLaVA-Bench (Wild) Heron-Bench Method Detail Conv Complex Avg. Detail Conv Complex Avg. Detail Conv Complex Avg. Results from Heron-Bench GPT-4V 88.0 84.7 97.5 90.1 89.9 93.1 99.1 94.1 83.3 77.5 78.3 79.7 Claude 3 Opus 67.9 70.4 95.3 77.9 87.4 73.4 94.6 85.1 74.5 68.4 77.7 73.6 Gemini Pro 65.3 83.0 75.4 74.6 61.7 84.2 84.2 76.7 55.6 64.3 64.0 61.3 LLaVA 1.6 7B 60.3 83.7 60.6 68.2 36.7 44.7 53.4 44.9 30.9 37.3 31.0 33.1 LLaVA 1.5 7B 71.1 83.7 69.6 74.8 49.2 48.7 54.7 50.8 42.4 45.9 35.5 41.3 Qwen-VL 7B 78.0 81.0 82.2 80.4 55.9 49.7 56.4 54.0 46.3 50.6 52.3 49.7 Japanese StableVLM 7B 18.9 54.8 24.1 32.6 26.0 24.8 29.2 26.7 25.2 51.2 37.8 38.1 EvoVLM-JP-v1 7B 61.0 75.7 71.0 69.2 49.6 65.5 54.2 56.4 50.3 44.4 40.5 45.1 Heron BLIP v1 7B 84.8 94.3 89.5 89.5 45.5 32.9 56.9 45.1 49.1 41.5 45.7 45.4 Heron GIT 7B 83.0 78.2 91.1 84.1 41.0 39.9 54.6 45.2 42.8 54.2 43.5 46.8 Our results LLaVA1.5-Swallow 8B 81.6 95.9 87.1 88.2 69.3 50.6 77.6 65.8 70.1 62.3 65.0 65.8 LLaVA1.5-Llama3 8B 82.8 90.6 87.2 86.9 60.5 38.7 71.5 56.9 68.7 60.1 56.1 61.6 LLaVA1.5-Phi3-mini 3.8B 77.3 89.8 83.7 83.6 66.7 36.5 66.8 56.7 61.6 61.2 48.5 57.1 LLaVA1.5-Phi3-medium 14B 84.7 89.6 86.2 86.8 77.0 68.2 77.0 74.1 62.0 56.1 54.1 57.4
Table 3: Main results. Our models trained with PDF data achieve high performance across all benchmarks except JA-LLaVA-Bench (COCO). Our LLaVA1.5-based models are named after their backbone LLMs (e.g., LLaVA1.5-Swallow). The best-performing open-source models are highlighted in bold.
@@ 5.1 Evaluation Metric
Heron-Bench follows the same score calculation method as LLaVA-Bench. First, GPT-4 (gpt-4-0125-previewOpenAI (2023) generates reference answers using the question’s context. Then, GPT-4 evaluates the answers of both the LMM and the reference answers using the LLM-as-a-judge approach Zheng et al. (2024). The final score is calculated as the ratio (%) of the average score of the LMM’s answers to the average score of GPT-4’s reference answers. A score of 100% indicates performance on par with GPT-4, while scores above 100% suggest that the LMM outperforms GPT-4.
@@ 5.2 Main Result
Table 3 presents the results of our model trained on PDF-derived data. Compared to existing models from the Heron-Bench paper, our LLaVA1.5-Swallow outperforms most open-source Japanese LMMs. It lags behind Heron BLIP v1 7B by only 1.3% on JA-LLaVA-Bench (COCO). On JA-LLaVA-Bench (Wild), it achieves a performance of 65.8%, surpassing the previous best of 56.4% by 9.4%. For Heron-Bench, it outperforms the previous best of 49.7% by a significant margin of 16.1%. Additionally, our LLaVA1.5-Llama3, LLaVA1.5-Phi3-mini, and LLaVA1.5-Phi3-medium also achieve higher performance than existing models. These results demonstrate that utilizing PDF data effectively enhances model performance.
@@ 5.3 Is Using PDF-derived Data Effective?
Table 4 presents the results of LLaVA1.5 training for each LLM across different stages. Overall, continual fine-tuning (CFT) on PDF-derived data is effective. Immediately after Stage 1 (Pretraining), all LLMs exhibit relatively low performance. After Stage 2 (Instruction tuning), Heron-Bench performance improves by approximately 20%–30%. After Stage 3 (CFT on PDF data), Heron-Bench performance further increases by at least 3.9% (Llama3) and up to 13.8% (Phi3-mini).
@@ 5.4 Which Japanese LLM Performs Best?
We trained LLaVA1.5 up to Stage 1 and 2 using the three commonly used Japanese LLMs introduced in §4.2 (Suzume, ELYZA, and Swallow). Table 5 presents the results. We found that Swallow achieved the best performance across all three benchmarks, and therefore, we selected Swallow as the Japanese LLM for this study.
@@ 5.5 What If We Use Raw Image-Text Pairs Without Generating Instruction Data?
Instead of generating instruction data from extracted image-text pairs, an LMM can be trained directly on them. We trained LLaVA1.5-Swallow with 50K PDF data using extracted image-text pairs. Table 6 presents the results.
@@ 5.6 Is Paired Text Effective for Generating Instruction Data?
When generating instruction data, we use paired text matched with images as context. However, the effectiveness of this paired text remains uncertain. Since the text extracted from PDFs is often imperfect and contains noise, its usefulness for training may be uncertain. To investigate this, we generate instruction data using different data sources and present the results in Table 7. We compare three cases: using only images, using images with paired text, and using images with PDF-style text.
@@
Table 7: Comparison of data sources for generating instruction data. LLaVA1.5-Swallow is used for these experiments.
@@ 6 Conclusion
We explore the use of Japanese PDF data to enhance LMM training and develop a fully automated pipeline for extracting image-text pairs. Our experiments show significant performance gains by incorporating PDF-derived data, with up to 13.8% improvement on Heron-Bench. Further analysis confirms the effectiveness of PDF-derived data across different model sizes and its potential to complement existing multimodal datasets. These findings provide valuable insights into leveraging PDF data for LMM training and highlight its promise as a multimodal resource. We hope this work encourages further research on leveraging PDF data for improving Japanese LMMs.
@@ Appendix A Prompts Used in Our Experiments
This section provides the prompts used in our experiments. Table A shows the prompt used to generate PDF-style text. Table B presents the prompt used to generate instruction data.
@@ Appendix B Qualitative Analysis
Figures A, B, and C present qualitative analyses. Each figure includes an image, a question, the reference answer from GPT-4, the response from LLaVA1.5-Swallow trained up to stages 1 and 2, and the response from LLaVA1.5-Swallow further trained with stage 3 (CFT on PDF). The results show that performance improves after training up to stage 3, demonstrating the effectiveness of CFT using PDF data.

### v2

@@ Abstract
Large Multimodal Models (LMMs) have demonstrated strong performance in English, but their effectiveness in Japanese remains limited due to the lack of high-quality training data. Current Japanese LMMs often rely on translated English datasets, restricting their ability to capture Japan-specific cultural knowledge. To address this, we explore the potential of Japanese PDF data as a training resource, an area that remains largely underutilized. We introduce a fully automated pipeline that leverages pretrained models to extract image-text pairs from PDFs through layout analysis, OCR, and vision-language pairing, removing the need for manual annotation. Additionally, we construct instruction data from extracted image-text pairs to enrich the training data. To evaluate the effectiveness of PDF-derived data, we train Japanese LMMs and assess their performance on the Japanese LMM Benchmark. Our results demonstrate substantial improvements, with performance gains ranging from 2.1% to 13.8% on Heron-Bench. Further analysis highlights the impact of PDF-derived data on various factors, such as model size and language models, reinforcing its value as a multimodal resource for Japanese LMMs.
@@
Jeonghun Baek Akiko Aizawa Kiyoharu Aizawa The University of Tokyo  National Institute of Informatics baek@hal.t.u-tokyo.ac.jp https://github.com/ku21fan/PDF-JLMM
@@ 1 Introduction
We train Japanese LMMs with the PDF-derived data and evaluate them on the Japanese LMM Benchmark. Our results show that PDF-derived data significantly improves performance, with gains of 2.1% to 13.8% on Heron-Bench Inoue et al. (2024). Furthermore, we conduct additional experiments to provide various insights into PDF-derived data’s effectiveness. Our main contributions are as follows:
@@
  • We demonstrate that PDF-derived data can significantly improve Japanese LMM performance, achieving 2.1% to 13.8% gains on Heron-Bench.
@@ 2.2 Japanese LMM
Some models, such as Qwen-VL Bai et al. (2023), achieve strong performance in Japanese without using a Japanese LLM, instead leveraging a multilingual LLM. VILA-jp Sasagawa et al. (2025) has further improved performance by utilizing interleaved data. MangaLMM Baek et al. (2025) specifically targets Japanese manga and demonstrates the potential of domain-specific Japanese LMMs. However, no existing work has leveraged PDF data to enhance Japanese LMMs. To achieve higher performance, we utilize PDF data in our approach. Recently, several benchmarks have been introduced for evaluating Japanese LMMs, such as JDocQA Onami et al. (2024), which focuses on Japanese document QA, and JMMMU Onohara et al. (2025), which considers Japanese culture.
@@ 4.1 Training Procedure
Dataset #Questions #Images JA-LLaVA-Bench (COCO) 90 30 JA-LLaVA-Bench (Wild) 60 24 Heron-Bench 103 21
Table 2: Details of Japanese LMM evaluation sets.
JA-LLaVA-Bench (COCO) JA-LLaVA-Bench (Wild) Heron-Bench Method Detail Conv Complex Avg. Detail Conv Complex Avg. Detail Conv Complex Avg. Results from Heron-Bench GPT-4V 88.0 84.7 97.5 90.1 89.9 93.1 99.1 94.1 83.3 77.5 78.3 79.7 Claude 3 Opus 67.9 70.4 95.3 77.9 87.4 73.4 94.6 85.1 74.5 68.4 77.7 73.6 Gemini Pro 65.3 83.0 75.4 74.6 61.7 84.2 84.2 76.7 55.6 64.3 64.0 61.3 LLaVA 1.6 7B 60.3 83.7 60.6 68.2 36.7 44.7 53.4 44.9 30.9 37.3 31.0 33.1 LLaVA 1.5 7B 71.1 83.7 69.6 74.8 49.2 48.7 54.7 50.8 42.4 45.9 35.5 41.3 Qwen-VL 7B 78.0 81.0 82.2 80.4 55.9 49.7 56.4 54.0 46.3 50.6 52.3 49.7 Japanese StableVLM 7B 18.9 54.8 24.1 32.6 26.0 24.8 29.2 26.7 25.2 51.2 37.8 38.1 EvoVLM-JP-v1 7B 61.0 75.7 71.0 69.2 49.6 65.5 54.2 56.4 50.3 44.4 40.5 45.1 Heron BLIP v1 7B 84.8 94.3 89.5 89.5 45.5 32.9 56.9 45.1 49.1 41.5 45.7 45.4 Heron GIT 7B 83.0 78.2 91.1 84.1 41.0 39.9 54.6 45.2 42.8 54.2 43.5 46.8 Our results PDF-JLMM 8B 81.6 95.9 87.1 88.2 69.3 50.6 77.6 65.8 70.1 62.3 65.0 65.8 LLaVA1.5-Llama3 8B 82.8 90.6 87.2 86.9 60.5 38.7 71.5 56.9 68.7 60.1 56.1 61.6 LLaVA1.5-Phi3-mini 3.8B 77.3 89.8 83.7 83.6 66.7 36.5 66.8 56.7 61.6 61.2 48.5 57.1 LLaVA1.5-Phi3-medium 14B 84.7 89.6 86.2 86.8 77.0 68.2 77.0 74.1 62.0 56.1 54.1 57.4
Table 3: Main results. Our models trained with PDF data achieve high performance across all benchmarks except JA-LLaVA-Bench (COCO). Our LLaVA1.5-based models are named after their backbone LLMs (e.g., LLaVA1.5-Llama3). PDF-JLMM refers to LLaVA1.5 with Swallow as the backbone LLM. The best-performing open-source models are highlighted in bold.
Training Data.  Table 1 presents the details of the training data. In Stage 1 (Pretraining) and Stage 2 (Instruction Tuning), we use the Japanese-translated version of the original LLaVA training data Inoue et al. (2024), translated using DeepL DeepL (2017). Specifically, Stage 1 uses 558K samples from LLaVA-Pretrain-JA Motors (2024a), Stage 2 uses 620K samples from LLaVA-v1.5-Instruct-620K-JA Motors (2024b), and Stage 3 uses 362K PDF-derived samples created in §3.
@@ 4.2 LLM Selection
Japanese LLM.  We train LLaVA1.5 using three well-known Llama3-8B-based Japanese LLMs: Suzume (suzume-llama-3-8B-japanese Devine (2024)), ELYZA (Llama-3-ELYZA-JP-8B Hirakawa et al. (2024)), and Swallow (Llama-3-Swallow-8B-Instruct-v0.1 Okazaki et al. (2024)) We build our main model by finetuning LLaVA1.5, using Swallow as the base Japanese LLM, with Japanese PDF data. We name this model PDF-JLMM (PDF-based Japanese Large Multimodal Model).
@@ 5.1 Evaluation Metric
Heron-Bench follows the same score calculation method as LLaVA-Bench. First, GPT-4 (gpt-4-0125-previewOpenAI (2023) generates reference answers using the question’s context. Then, GPT-4 evaluates both LMM’s answers and the reference answers using the LLM-as-a-judge approach Zheng et al. (2024). The final score is calculated as the ratio (%) of the average score of the LMM’s answers to the average score of GPT-4’s reference answers. A score of 100% indicates performance on par with GPT-4, while scores above 100% suggest that the LMM outperforms GPT-4.
@@ 5.2 Main Result
Table 3 presents the results of our model trained on PDF-derived data. Compared to existing models from the Heron-Bench paper, our PDF-JLMM outperforms most open-source Japanese LMMs. It lags behind Heron BLIP v1 7B by only 1.3% on JA-LLaVA-Bench (COCO). On JA-LLaVA-Bench (Wild), it achieves a performance of 65.8%, surpassing the previous best of 56.4% by 9.4%. For Heron-Bench, it outperforms the previous best of 49.7% by a significant margin of 16.1%. Additionally, our LLaVA1.5-Llama3, LLaVA1.5-Phi3-mini, and LLaVA1.5-Phi3-medium also achieve higher performance than existing models. These results demonstrate that utilizing PDF data effectively enhances model performance.
@@ 5.3 Is Using PDF-derived Data Effective?
Table 4 presents the results of LLaVA1.5 training for each LLM across different stages. Overall, continual fine-tuning (CFT) on PDF-derived data is effective. Immediately after Stage 1 (Pretraining), all LLMs exhibit relatively low performance. After Stage 2 (Instruction tuning), Heron-Bench performance improves by approximately 20%–30%. After Stage 3 (CFT on PDF data), Heron-Bench performance further increases by at least 2.1% (Phi3-medium) and up to 13.8% (Phi3-mini).
@@ 5.4 Which Japanese LLM Performs Best?
We train LLaVA1.5 up to Stage 1 and 2 using the three commonly used Japanese LLMs introduced in §4.2 (Suzume, ELYZA, and Swallow). Table 5 presents the results. We find that Swallow achieves the best performance across all three benchmarks, and therefore, we select Swallow as the Japanese LLM for this study.
@@ 5.5 What If We Use Raw Image-Text Pairs Without Generating Instruction Data?
Instead of generating instruction data from extracted image-text pairs, an LMM can be trained directly on them. We train PDF-JLMM with 50K PDF data using extracted image-text pairs. Table 6 presents the results.
@@ 5.6 Is Paired Text Effective for Generating Instruction Data?
When generating instruction data, we use paired text matched with images as context. However, the effectiveness of this paired text remains uncertain. Since the text extracted from PDFs is often imperfect and contains noise, its usefulness for training may be uncertain. To investigate this, we generate instruction data using different data sources derived from 50K PDFs and present the results in Table 7. We compare three cases: using only images, images with paired text, and images with PDF-style text.
@@
Table 7: Comparison of data sources for generating instruction data. PDF-JLMM is used in experiments.
5.7 Comparison with the Translated Datasets
We compare the PDF-derived data with Japanese-translated versions of existing English instruction data. The 620K instruction data (LLaVA-v1.5-Instruct-620K-JA) used in Stage 2 is considered a different Japanese multimodal dataset, constructed by translating English instruction data. We compare it with our Instruct-from-200K PDF dataset (362K instruction data). Instead of using LLaVA-v1.5-Instruct-620K-JA in Stage 2, we train the model with our Instruct-from-200K PDF dataset. Table 8 shows the results. LLaVA1.5-Swallow 8B trained with the PDF-derived data outperforms the one trained with LLaVA-v1.5-Instruct-620K-JA, even with fewer samples. This result suggests that the performance gains are not merely due to large-scale instruction tuning, but rather from leveraging instruction data constructed from Japanese PDFs containing culturally grounded content.
Data source for Stage 2 L-COCO L-Wild Heron LLaVA-v1.5-Instruct-620K-JA 84.0 59.8 54.7 Instruct-from-200K PDF 88.1 72.7 70.0
Table 8: Comparison with the translated data, LLaVA-v1.5-Instruct-620K-JA.
We also compare the PDF-derived data with other datasets that we translated into Japanese. Specifically, we translated Vision-Flan Xu et al. (2024) (186K instruction data) and Image-Textualization Pi et al. (2024) (99.6K instruction data)—subsets of LLaVA OneVision Li et al. (2024)—into Japanese using GPT-4o-mini. For translation, we use the prompt in Table C. We use these datasets for CFT in Stage 3. Table 9 shows the results. The results show that simply translating effective English data into Japanese may even lead to performance degradation on Japanese benchmarks. This underscores the importance of using Japanese PDF data that contains authentic, culturally grounded content.
Our approach differs from translated English datasets in that we extract images containing authentic Japanese content directly from Japanese PDFs. For example, as shown in Figure A of the qualitative analysis in the supplementary materials, the model trained only up to Stage 2 (before incorporating our PDF-derived data) describes “cherry blossoms”—a culturally symbolic flower in Japan—as “white flowers”, whereas the model further trained on PDF-derived data correctly refers to it as “cherry blossoms”. We believe this illustrates a key distinction from translated data, as our method helps the model learn cultural concepts grounded in native Japanese contexts rather than relying on translations.
Data source for Stage 3 L-COCO L-Wild Heron 100K PDF (181K inst.) 88.2 65.8 65.8 Vision-Flan (VF) 37.8 29.0 40.2 Image-Textualization (IT) 81.6 43.3 46.3 VF + IT 56.7 35.5 39.9
Table 9: Comparison with the translated subset of LLaVA-OneVision. LLaVA1.5-Swallow is used in experiments.
@@ 6 Conclusion
We explore the use of Japanese PDF data to enhance LMM training and develop a fully automated pipeline for extracting image-text pairs. Our experiments show significant performance gains by incorporating PDF-derived data, with up to 13.8% improvement on Heron-Bench. Further analysis confirms the effectiveness of PDF-derived data across different model sizes and its potential to complement existing multimodal datasets. These findings provide valuable insights into leveraging PDF data for LMM training and highlight its promise as a multimodal resource. While our focus is on Japanese, we believe this approach is applicable to other languages as well, and we hope it fosters further research toward improving LMMs across diverse languages.
@@ Appendix A Prompts Used in Our Experiments
This section provides the prompts used in our experiments. Table A, Table B, and Table C show the prompts used to generate PDF-style text, instruction data, and Japanese translations of English datasets, respectively.
@@
[…]
Table C: Prompt for translating English datasets into Japanese datasets. Vision-Flan Xu et al. (2024) and Image-Textualization Pi et al. (2024), subsets of LLaVA OneVision Li et al. (2024), are translated into Japanese using GPT-4o-mini.
@@ Appendix B Qualitative Analysis
Figures A, B, and C present qualitative analyses. Each figure includes an image, a question, the reference answer from GPT-4, the response from LLaVA1.5-Swallow trained up to stages 1 and 2, and the response from LLaVA1.5-Swallow further trained with stage 3 (CFT on PDF). The results show that performance improves after training up to stage 3, demonstrating the effectiveness of CFT using PDF-derived data.